If 3 apples cost 48 cents how many dozen apples can be bought for 3.84....please show how you got answer
Answer by harpazo(655)

3/a = 0.48/3.84

0.48a = 3(3.84)
0.48a = 11.52
a = 11.52/0.48
a = 24
Answer is 24 apples.